OpenStudy (anonymous):

Find the slope equation and the line given the characteristics slope -6 and y intercept (0,4) y=

OpenStudy (anonymous):

the point slope formula is y-y1 = m(x-x1) x1 and y1 are your coordinates and m is the slope. so you can plug the numbers in and solve for y (:

OpenStudy (campbell_st):

use the point slope formula \[y - y _{1} = m(x - x _{1})\] y - 4 = -6(x - 0) expand and simplify

OpenStudy (anonymous):

thanx c an j